OpenStudy (anonymous):

Perform the indicated operation. (-9 + 2i) - (-12 + 4i) =

OpenStudy (callisto):

(-9 + 2i) - (-12 + 4i) = (-9 + 2i +12 - 4i) = 3 - 2i